# Antidote: Post-fine-tuning Safety Alignment for Large Language Models against Harmful Fine-tuning

### v2

@@ Appendix A Experiment Setup
For different fine-tuning tasks, we accordingly construct the triplet of Instruction/Input/Response. For example, for SST2 tasks, the instruction is "Analyze the sentiment of the input, and respond only positive or negative", the input is the according sentence in SST2 dataset, and the response is the according label of the sentence, i.e., "positive" or "negative". For SST2, AGNEWS, and GSM8K, we measure the finetune accuracy by counting the correct samples out of all the testing samples. A sample is counted as correct for SST2 and AGNEWS if the model gives the correct classification answer. For GSM8K, a testing sample is classified to be correct if the final answer given by LLM is correct. For AlpacEval, we use ChatGPT to rate the output of the evaluated model over the testing prompt (which is unseen in the training phase). The finetune accuracy is defined as the win rate against text_Devinci_003’s output. The measurement method is consistent with previous work (Huang, Hu, and Liu 2024; Huang et al. 2024).
Implementation of Baselines and Their Idea
@@
  • Vaccine (alignment-stage solution). For Vaccine (Huang, Hu, and Liu 2024), we use Vaccine to align the pre-trained model on the alignment dataset. Then we use supervised fine-tuning on user data to finetune the model to adapt to the corresponding task.
  • RepNoise (alignment-stage solution). For RepNoise (Rosati et al. 2024a), we use RepNoise to align the pre-trained model on the alignment dataset/harmful dataset. Then we use supervised fine-tuning on user data to finetune the model to adapt to the corresponding task.
  • Lisa (fine-tuning-stage solution). For Lisa (Huang et al. 2024), we use SFT to align the pre-trained model on the alignment dataset. Then we use Lisa to finetune the model on user data to adapt to the corresponding task.
  • LDIFS (fine-tuning-stage solution). For LDIFS (Mukhoti et al. 2023) , we use SFT to align the pre-trained model on the alignment dataset. Then we use LDIFS to finetune the model on user data to adapt to the corresponding task.
@@
  • Vaccine (alignment-stage solution). The core idea of Vaccine (Huang, Hu, and Liu 2024) is to add perturbation to the hidden embedding in the alignment stage, such that the produced embedding is able to resist the real harmful perturbation in the fine-tuning stage (i.e., to vaccinate the model). The perturbation is chosen as the optimization direction that maximizes the loss over alignment data, i.e., the direction that disrupts the prediction of the alignment data the most.
  • RepNoise (alignment-stage solution). The core contribution of RepNoise (Rosati et al. 2024a) is the representation noise. Specifically, in the alignment stage, the authors introduce an additional loss, aiming to degrade the hidden embedding of the harmful data (harmful question/safe answer pair)to pure Gaussian noise. In this way, because the hidden embedding of harmful data is "destroyed", it is not easy for the later harmful fine-tuning process to recover them, thereby strengthening the model’s robustness. RepNoise assumes both the availability of harmful data (i.e., realignment data for Antidote) and the alignment data (i.e., harmful question/harmful answer pair).
@@
  • LDIFS (fine-tuning-stage solution). For LDIFS (Mukhoti et al. 2023) , the idea is to make the hidden embedding of the fine-tuning data of the current model closer to the embedding of the original aligned model. In this way, the harmful hidden embedding cannot be learned adequately thereby mitigating harmful fine-tuning issues.

### v3

@@ A.1 Detailed Setup
For different fine-tuning tasks, we accordingly construct the triplet of Instruction/Input/Response. For example, for SST2 tasks, the instruction is "Analyze the sentiment of the input, and respond only positive or negative", the input is the according sentence in SST2 dataset, and the response is the according label of the sentence, i.e., "positive" or "negative". For SST2, AGNEWS, and GSM8K, we measure the finetune accuracy by counting the correct samples out of all the testing samples. A sample is counted as correct for SST2 and AGNEWS if the model gives the correct classification answer. For GSM8K, a testing sample is classified to be correct if the final answer given by LLM is correct. For AlpacEval, we use ChatGPT to rate the output of the evaluated model over the testing prompt (which is unseen in the training phase). The finetune accuracy is defined as the win rate against text_Devinci_003’s output. The measurement method is consistent with previous work (Huang et al., 2024d, b).
A.2 Implementation of Baselines and Their Idea
@@
  • Vaccine (alignment-stage solution). For Vaccine (Huang et al., 2024d), we use Vaccine to align the pre-trained model on the alignment dataset. Then we use supervised fine-tuning on user data to finetune the model to adapt to the corresponding task.
  • RepNoise (alignment-stage solution). For RepNoise (Rosati et al., 2024a), we use RepNoise to align the pre-trained model on the alignment dataset/harmful dataset. Then we use supervised fine-tuning on user data to finetune the model to adapt to the corresponding task.
  • Lisa (fine-tuning-stage solution). For Lisa (Huang et al., 2024b), we use SFT to align the pre-trained model on the alignment dataset. Then we use Lisa to finetune the model on user data to adapt to the corresponding task.
  • LDIFS (fine-tuning-stage solution). For LDIFS (Mukhoti et al., 2023) , we use SFT to align the pre-trained model on the alignment dataset. Then we use LDIFS to finetune the model on user data to adapt to the corresponding task.
@@
  • Vaccine (alignment-stage solution). The core idea of Vaccine (Huang et al., 2024d) is to add perturbation to the hidden embedding in the alignment stage, such that the produced embedding is able to resist the real harmful perturbation in the fine-tuning stage (i.e., to vaccinate the model). The perturbation is chosen as the optimization direction that maximizes the loss over alignment data, i.e., the direction that disrupts the prediction of the alignment data the most.
  • RepNoise (alignment-stage solution). The core contribution of RepNoise (Rosati et al., 2024a) is the representation noise. Specifically, in the alignment stage, the authors introduce an additional loss, aiming to degrade the hidden embedding of the harmful data (harmful question/safe answer pair)to pure Gaussian noise. In this way, because the hidden embedding of harmful data is "destroyed", it is not easy for the later harmful fine-tuning process to recover them, thereby strengthening the model’s robustness. RepNoise assumes both the availability of harmful data (i.e., realignment data for Antidote) and the alignment data (i.e., harmful question/harmful answer pair).
@@
  • LDIFS (fine-tuning-stage solution). For LDIFS (Mukhoti et al., 2023) , the idea is to make the hidden embedding of the fine-tuning data of the current model closer to the embedding of the original aligned model. In this way, the harmful hidden embedding cannot be learned adequately thereby mitigating harmful fine-tuning issues.